Thane Company is interested in establishing the relationship between electricity costs and machine hours. Data have been collected and a regression analysis prepared using Excel. The monthly data and the regression output follow: Month Machine Hours Electricity Costs January 2,500 $ 19,200 February 2,900 $ 22,600 March 1,900 $ 14,300 April 3,100 $ 24,600 May 3,800 $ 29,050 June 3,300 $ 23,600 July 4,100 $ 25,550 August 3,500 $ 23,550 September 2,000 $ 17,100 October 3,700 $ 27,600 November 4,700 $ 32,600 December 4,200 $ 28,550 Summary Output Regression Statistics Multiple R .956 R Square .914 Adjusted R2 .905 Standard Error 1,615.39 Observations 12.00 Coefficients Standard Error t Stat P-value Lower 95% Upper 95% Intercept 4,974.43 1,907.43 2.61 0.03 724.41 9,224.45 Machine Hours 5.76 0.56 10.30 0.00 4.51 7.00

If the controller uses the high-low method to estimate costs, the fixed cost portion of the cost equation for electricity cost is: (Round intermediate calculations to 2 decimal places.) $1,615.39. $1,862.00. $4,974.43. $24,138.00.
Answer #1
Machine hours Electricity Costs
High-level 4700 32600
Low-level 1900 14300
Difference 2800 18300
Unit variable cost = 18300/2800 = $6.54
Fixed cost portion = 32600-(4700*6.54)= $1862.00
Option 2 $1,862.00 is correct